what is th slope of the line that passes through (6,-7) and (4,-8)
Answer by jim_thompson5910(35256)

Two generic points that define a straight line are (x1, y1) and (x2, y2)

In this case, (x1, y1) = (6, -7), which means x1=6 and y1=-7

and (x2, y2) = (4, -8), which means x2=4 and y2=-8

So x1=6, y1=-7, x2=4, and y2=-8.

-------------------


Now use the slope formula


m = (y2-y1)/(x2-x1)


m = (-8-(-7))/(4-6)


m = (-8+7)/(4-6)


m = (-1)/(-2)


m = 1/2


So the slope of the line that passes through the two points is 1/2